Write The rational numbers that are equal to their reciprocals.

Write 
 The rational numbers that are equal to their reciprocals. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The rational numbers that are equal to their reciprocals are 1 and -1.
Reason:
1=1/1
Reciprocal of 1 = 1/1 = 1 Similarly, Reciprocal of -1 = - 1

Thanks